a particle is dropped from a height h and at the same time another particle is projected vertically up frm the ground . they meet when the first has descended a height of h/3 .find ratio of their velocities at this instant

Grade:11

1 Answers

bandana
5 years ago
at h/3 they meet
falling body reached – h/3
body thrown upward – h-h/3 = (2h)/3
they are thrown at same time, therefore by equating time ( time= distance/speed)
v1/v2 = 1/2